# TC-Light: Temporally Coherent Generative Rendering for Realistic World Transfer

### v2

@@ Abstract
Illumination and texture editing are critical dimensions for world-to-world transfer, which is valuable for applications including sim2real and real2real visual data scaling up for embodied AI. Existing techniques generatively re-render the input video to realize the transfer, such as video relighting models and conditioned world generation models. Nevertheless, these models are predominantly limited to the domain of training data (e.g., portrait) or fall into the bottleneck of temporal consistency and computation efficiency, especially when the input video involves complex dynamics and long durations. In this paper, we propose TC-Light, a novel generative renderer to overcome these problems. Starting from the video preliminarily relighted by an inflated video relighting model, it optimizes appearance embedding in the first stage to align global illumination. Then it optimizes the proposed canonical video representation, i.e., Unique Video Tensor (UVT), to align fine-grained texture and lighting in the second stage. To comprehensively evaluate performance, we also establish a long and highly dynamic video benchmark. Extensive experiments show that our method enables physically plausible re-rendering results with superior temporal coherence and low computation cost. The code and video demos are available at https://dekuliutesla.github.io/tclight/.
@@ 1 Introduction
Lighting and its interaction with both real and synthetic environments fundamentally shapes how humans—and embodied agents—perceive the world. The ability to re-render the illumination and texture (or so-called relighting) of camera-captured image sequences, especially in complex, highly dynamic scenes, is critically valuable for various world-to-world transfer use cases. For instance, by re-rendering the CG-simulated or realistic video data used to train embodied agents, it can bridge the sim-to-real gap and enable real-to-real transfer, thus unlocking access to massive high-quality data that is essential for stepping towards embodied intelligence.
Despite its importance, the video illumination re-rendering remains a highly challenging problem, particularly when camera motion is highly dynamic and foreground objects frequently enter and exit scenes, which are common cases for embodied agents, as shown in Fig. 1. Most existing generative relighting techniques zhang2025scaling ; li2022physically ; jin2024neural_gaffer ; ren2024relightful ; kim2024switchlight ; wang2023sunstage are tailored for static images. As shown in Sec. 4.2, naively inflating them to a video model with existing zero-shot strategies struggles to balance the consistency and quality. Moreover, the considerable training cost and scarcity of video lighting datasets hinder fine-tuning a pretrained model for this task. Besides, though generative video relighting and world generation models are emerging, they are either restricted on domain of training data zhang2021neural ; choi2024personalized ; cai2024real ; alhaija2025cosmos or burdened by considerable computation overhead zhou2025light ; fang2025relightvid on long video, as validated in Sec. 4.2.
@@
To address the limitations outlined above, we propose TC-Light. We utilize the SOTA image relighting model IC-Light zhang2025scaling as the baseline, and inflate it to a video model in a zero-shot manner with incorporation of our proposed decayed multi-axis denoising. It provides a preliminary video relighting result. The core innovation of TC-Light lies in a two-stage post-optimization framework that substantially improves temporal consistency. The first stage introduces per-frame appearance embedding to compensate for exposure discrepancy. It is optimized with photometric loss against the preliminarily relighted video and a flow-based loss between adjacent frames. This enforces global illumination consistency and facilitates consequent optimization. The second stage compresses the output to a canonical representation, i.e., Unique Video Tensor (UVT), according to priors including optical flow and depth of the source video. UVT is then optimized by minimizing the warping error across decompressed frames while aligning the content with the first stage result. As shown in Tab. 2, our optimization procedure is extremely efficient and introduces minimal VRAM overhead.
To comprehensively assess the effectiveness of our model, we introduce a challenging benchmark tailored for complex and highly dynamic scenes. It comprises 58 videos of averagely 256 frames per clip, spanning both indoor and outdoor environments, realistic and synthetic settings, and a wide range of lighting and weather conditions. Extensive experiments demonstrate that our method achieves high-quality, temporally consistent video relighting while maintaining low computational overhead, highlighting its strong potential for downstream applications such as embodied AI. Our main contributions are as follows:
@@ 3.1 Preliminaries
Task Setting. As shown in Fig. 2, we take RGB video as input. The axes of the video space-time volume are denoted by $(x,y,t)$, where $xy$ planes correspond to video frames and $yt$ planes are defined as spatiotemporal slices cohen2024slicedit . Since the camera motion is highly dynamic, the target illumination can no longer be simply appointed by a static image or an HDR environment map. Due to superior flexibility and operability, we use textual prompts as the control signal and relight the entire frame.
@@
where $\{\alpha_{\tau}\}$ is a monotonically decreasing noise schedule. The reverse process begins from pure noise $z_{T}\sim\mathcal{N}(0,\textbf{{I}})$. With guidance from control signal (image, text, depth, etc) $c$, the trained UNet ronneberger2015u $\epsilon_{\theta}$ estimates the noise direction and progressively removes the noise from $z_{T}$. After the final denoising step, the estimated clean latent $\hat{z_{0}}$ is decoded by $\mathcal{D}(\cdot)$ to obtain the generated image $\hat{x_{0}}=\mathcal{D}(\hat{z_{0}})$, which approximates the training distribution.
Figure 2: TC-Light overview. Given the source video and text prompt $p$, the model tokenizes input latents in $xy$ plane and $yt$ plane seperately. The predicted noises are combined together for denoising (cf. Sec. 3.2). Its output then undergoes two-stage optimization to enhance temporal consistency of illumination and texture, which are respectively detailed in Sec. 3.3.1 and Sec. 3.3.2.
@@ 3.2 Lifting Image Diffusion Model to Video Space
For (ii), we propose decayed multi-axis denoising. Similar to Slicedit cohen2024slicedit , the denoiser has two components with shared weights $\epsilon^{xy}_{\theta}(\cdot,p)$ that tokenizes each frame and merges tokens from local temporal slots, while $\epsilon^{yt}_{\theta}(\cdot,``\,")$ tokenizes the $yt$ planes (cf. Sec. 3.1) and merges tokens from local image width slot. Note that $\epsilon^{xy}_{\theta}$ conditions on target prompt $p$, while $\epsilon^{yt}_{\theta}$ takes empty prompt ``"``"``\,"` ` " as input (making the denoiser unconditional). The noises separately predicted by two parts according to the same input latents are combined together cohen2024slicedit
$\epsilon^{V}_{\theta}(\cdot,p)=\sqrt{\gamma}\epsilon^{xy}_{\theta}(\cdot,p)+\sqrt{1-\gamma}\epsilon^{yt}_{\theta}(\cdot,``\,"),$ (2)
where hyperparameter $\gamma\in[0,1]$ balances effect from $\epsilon^{yt}_{\theta}$. However, the unconditional $\epsilon^{yt}_{\theta}$ would overly biases texture and lighting toward the source video, and therefore lead to unnatural relighting results, as validated in Fig. 3 and Fig. 4. To alleviate this problem, we replace $\gamma$ with a timestep-dependent $\gamma_{\tau}$ that exponentially decays during denoising. To further align predicted noise from $\epsilon^{yt}_{\theta}$ to that of $\epsilon^{xy}_{\theta}$, we use Adaptive Instance Normalization (AIN) huang2017arbitrary to adjust feature statistics
$\epsilon_{\theta}^{V}(\cdot,p)=\sqrt{\gamma_{\tau}}\epsilon_{\theta}^{xy}(\cdot,p)+\sqrt{1-\gamma_{\tau}}\widehat{\epsilon_{\theta}^{yt}}(\cdot,``\,"),$ (3)
@@
where $\mu_{*}$ and $\sigma_{*}$ are the channel-wise mean and standard deviation of each frame. This design preserves motion guidance from the source video while reducing unwanted texture and lighting bias, as validated by ablation studies in Sec. 4.3. The output denoised video is denoted as $\{I_{t}\}$.
@@ 3.3 Post Optimization for Temporal Consistency
Although the video diffusion extension in Sec. 3.2 has introduced spatial-temporal awareness and motion prior from source video, noticeable illumination and texture flicker persist. To efficiently remove these artifacts, we introduce a two-stage post-optimization framework, as illustrated in parts (b) and (c) of Fig. 2.
@@ 3.3.1 Stage I: Exposure Alignment
As shown in part (b) of Fig. 2, the first stage introduces a per-frame appearance embedding $E_{t}$ to compensate for exposure misalignment between adjacent frames. Inspired by kerbl2024hierarchical , we model $E_{t}$ as a $3\times 4$ affine transformation matrix, initialized to the identity and optimized via Adam 2015-kingma . Its supervision combines a photometric term with a flow-warp alignment term using hyperparameter $\lambda_{e}$
@@
Here, $\beta$ is a constant scaling factor, $\xi_{flow}$ and $\xi_{rgb}$ are thresholds set from the statistics of error map $E_{flow}$ and $E_{rgb}$. This soft mask is also applied in the second stage of optimization. As shown in Tab. 3, soft masking outperforms the hard one in both temporal consistency and prompt alignment.
@@ 3.3.2 Stage II: Optimization over Unique Video Tensor
In the second stage, we refine illumination and texture details. Compared with vanilla video, its canonical representation can incorporate spatial-temporal priors and facilitate consistency ouyang2024codef ; shin2024enhancing . But popular NeRF or 3DGS are too complex and costly for learning (cf. Sec. 2.2). Instead, we compress the video to a one-dimensional RGB vector of shape $(N,3)$, as shown in part (c) of Fig. 2. Specifically, we define a $d$-dimensional index $\kappa(x,y,t)$ for each pixel based on priors extracted from the source video. An example index could be $[22,127,0,255]$, where the first element is flow ID (grouping pixels with shared optical flow), and the rest are 8-bit quantized RGB values. It is also allowed to extend this 4-element index to more elements with voxel coordinate (from depth projection) or any other cues that indicate spatial–temporal similarity and locality. All pixels with identical $\kappa$ are gathered via averaging to form one element of the one-dimensional vector, where $N$ is the number of unique $\kappa$. Take the source video $\{I_{t}^{in}\}$ as an example, the gathering and scattering operations are formulated as
@@
where $\mathbf{U}$ is referred to as the Unique Video Tensor (UVT). With an appropriate definition of $\kappa$, the scattered $\mathbb{I}_{t}^{in}\left(x,y\right)$ reconstructs the original $I_{t}^{in}\left(x,y\right)$ with minimal information loss, as validated in Tab. 4. For relighting, the ideal edited video frames must preserve consistent motion and intrinsic image details with the source; thus, they share the same index tensor $\kappa$ for UVT representation. Accordingly, we compress the first-stage output $\tilde{I}_{t}\left(x,y\right)$ into $\mathbf{\tilde{U}}$ via Eq. 8, which then serves as the primary optimization target. This formulation not only facilitates optimization but also naturally embeds spatial-temporal similarity priors (cf. Sec. 4.3). With CUDA parallelism, the gathering and scattering process can be performed instantly. The optimization of $\tilde{\mathbf{U}}$ is supervised by
@@
where $\tilde{\mathbb{I}}_{t}\left(x,y\right)=\tilde{\mathbf{U}}\left(\kappa(x,y,t)\right)$, and $\lambda_{tv}$ and $\lambda_{u}\in[0,1]$ balance the loss terms. The total variation loss $\mathcal{L}_{tv}$ suppresses noise. Notably, Eq. 9 applies SSIM loss instead of photometric loss. This leaves space to fine-grained appearance and illumination adjustment without altering image structure. Finally, the optimized $\widehat{\mathbf{U}}$ is used to reconstruct $\widehat{\mathbb{I}}_{t}(x,y)$ according to Eq. 8 as the final output.
@@ 4.1 Experiment Setting
Dataset. To comprehensively evaluate the generation capability, we collect video clips with high motion dynamic and broad scene diversity. This benchmark, as detailed in Tab. 1, covers scenarios like autonomous driving, robot manipulation, and navigation, as well as drone flight. It includes data collected from synthetic and realistic environments under various weather conditions. Each clip is a long video with high dynamics and on average 256 frames, making it extremely challenging. To obtain edit prompts, we use some prompts from zhang2025scaling and generate others using COSMOS agarwal2025cosmos .
@@
Metrics. Following prior works qi2023fatezero ; li2024vidtome ; kara2024rave ; wang2025videodirector , we assess the relighting performance along following four dimensions: (i) Temporal consistency is quantified via motion smoothness (Motion-S) huang2024vbench and structural warping error (Warp-SSIM). Motion-S evaluates the continuity and physical plausibility of motion in the edited sequence, whereas Warp-SSIM computes the SSIM between a frame and its warped neighbors using flow from RAFT teed2020raft . (ii) Textual alignment is measured by average CLIP embedding similarity between the text prompt and all edited frames (CLIP-T). (iii) User preference is evaluated by a study on 19 randomly selected videos and 65 valid submissions collected. Participants choose their preferred relighting results among our method and established baselines, from which we derive the Bradley–Terry preference rate (User-PF) bradley1952rank . Additional details are included in the Appendix. (iv) Computation efficiency is reported in terms of runtime speed (FPS) and peak GPU memory consumption (VRAM) during editing. All experiments are conducted on a 40GB A100 GPU. Additionally, to appraise the reconstruction quality of UVT, we report average PSNR, SSIM, and LPIPS between the original and reconstructed frames.

### v3

@@ Abstract
Illumination and texture rerendering are critical dimensions for world-to-world transfer, which is valuable for applications including sim2real and real2real visual data scaling up for embodied AI. Existing techniques generatively re-render the input video to realize the transfer, such as video relighting models and conditioned world generation models. Nevertheless, these models are predominantly limited to the domain of training data (e.g., portrait) or fall into the bottleneck of temporal consistency and computation efficiency, especially when the input video involves complex dynamics and long durations. In this paper, we propose TC-Light, a novel paradigm characterized by the proposed two-stage post optimization mechanism. Starting from the video preliminarily relighted by an inflated video relighting model, it optimizes appearance embedding in the first stage to align global illumination. Then it optimizes the proposed canonical video representation, i.e., Unique Video Tensor (UVT), to align fine-grained texture and lighting in the second stage. To comprehensively evaluate performance, we also establish a long and highly dynamic video benchmark. Extensive experiments show that our method enables physically plausible re-rendering results with superior temporal coherence and low computation cost. The code and video demos are available at our Project Page.
@@ 1 Introduction
Lighting and its interaction with both real and synthetic environments fundamentally shapes how humans—and embodied agents—perceive the world. The ability to re-render the illumination and texture (or so-called relighting) of captured image sequences, especially in complex, highly dynamic scenes, is critically valuable for various world-to-world transfer use cases like filmmaking (Richardt_Stoll_Dodgson_Seidel_Theobalt_2012) and augmented reality (li2022physically). Crucially, by re-rendering the CG-simulated or realistic video data used to train embodied agents, it can bridge the sim-to-real gap and enable real-to-real transfer, thus unlocking access to massive high-quality data that is essential for stepping towards embodied intelligence.
Despite its importance, the video illumination and texture rerendering remains a highly challenging problem, particularly when camera motion is highly dynamic and foreground objects frequently enter and exit scenes, as shown in Fig.˜1. Most existing generative relighting techniques zhang2025scaling; li2022physically; jin2024neural_gaffer; ren2024relightful; kim2024switchlight; wang2023sunstage are tailored for static images. As shown in Sec.˜4.2, naively inflating them to a video model with existing zero-shot strategies struggles to balance the consistency and quality. Moreover, the considerable training cost and scarcity of video lighting datasets hinder fine-tuning a pretrained model for this task. Besides, though generative video relighting and world generation models are emerging, they are either restricted on domain of training data zhang2021neural; choi2024personalized; cai2024real; alhaija2025cosmos or burdened by considerable computation overhead zhou2025light; fang2025relightvid on long video, as validated in Sec.˜4.2.
@@
To address the limitations outlined above, we propose TC-Light. We utilize the SOTA image relighting model IC-Light zhang2025scaling as the baseline, and inflate it to a video model in a zero-shot manner with our decayed multi-axis denoising model, which is distinguished by the proposed Decayed Noise Weighting and Noise Statistic Alignment. It provides a preliminary video relighting result. The core innovation of TC-Light lies in a two-stage post-optimization framework that substantially improves temporal consistency. The first stage introduces per-frame appearance embedding to compensate for exposure discrepancy. It is optimized with photometric loss against the preliminarily relighted video and a flow-based loss between adjacent frames. This enforces global illumination consistency and facilitates consequent optimization. The second stage compresses the output to a canonical representation, i.e., Unique Video Tensor (UVT), according to priors including optical flow and depth of the source video. UVT is then optimized by minimizing the warping error across decompressed frames while aligning the content with the first stage result. As shown in Tab.˜2, our optimization procedure is extremely efficient and introduces minimal VRAM overhead.
To comprehensively assess the effectiveness of our model, we introduce a challenging benchmark tailored for complex and highly dynamic scenes. It comprises 58 videos of averagely 256 frames per clip, spanning both indoor and outdoor environments, realistic and synthetic settings, and a wide range of lighting and weather conditions. Extensive experiments demonstrate that our method achieves high-quality, temporally consistent video relighting while maintaining low computational overhead, highlighting its great potential for downstream applications such as embodied AI. Our main contributions are as follows:
@@ 3 Method
In this section, we first introduce the task setting and preliminary knowledge about latent diffusion models in Sec.˜3.1. Sec.˜3.2 further illustrates how our proposed Decayed Noise Weighting and Noise Statistics Alignment helps effectively lift the image relighting model to video space. Sec.˜3.3 details how our key innovation, i.e., the two-stage temporal consistency optimization strategy, helps align overall illumination and texture appearance.
@@ 3.1 Preliminaries
Task Setting. As shown in Fig.˜2, we take RGB video as input. The axes of the video space-time volume are denoted by $(x,y,t)$, where $xy$ planes correspond to video frames and $yt$ planes are defined as spatiotemporal slices cohen2024slicedit. Since the camera motion is highly dynamic, the target illumination can no longer be simply appointed by a static image or an HDR environment map. Due to superior flexibility and operability, we use textual prompts as the control signal and relight the entire frame.
@@
[…]
Figure 2: TC-Light overview. Given the source video and text prompt $p$, the model tokenizes input latents in xyxy plane and ytyt plane separately. The predicted noises are adaptively combined together for denoising (cf. Sec.˜3.2). Its output then undergoes two-stage optimization to enhance temporal consistency of illumination and texture, which are respectively detailed in Sec.˜3.3.1 and Sec.˜3.3.2.
@@ 3.2 Lifting Image Diffusion Model to Video Space
[…]
where hyperparameter $\gamma\in[0,1]$ balances effect from $\epsilon^{yt}_{\theta}$. However, the unconditional $\epsilon^{yt}_{\theta}$ would overly biases texture and lighting toward the source video, and therefore lead to unnatural relighting results, as validated in Fig.˜3 and Fig.˜4. To alleviate this problem, we introduce Decayed Noise Weighting, which replaces $\gamma$ with a timestep-dependent $\gamma_{\tau}$ that exponentially decays during denoising. To further align predicted noise from $\epsilon^{yt}_{\theta}$ to that of $\epsilon^{xy}_{\theta}$, we use Adaptive Instance Normalization (AIN) huang2017arbitrary to align noise statistics
[…]
@@
where $\mu_{*}$ and $\sigma_{*}$ are the channel-wise mean and standard deviation of each frame. This design preserves motion guidance from the source video while reducing unwanted texture and lighting bias, as validated by ablation studies in Sec.˜4.3. The output denoised video is denoted as $\{I_{t}\}$.
@@ 3.3 Post Optimization for Temporal Consistency
Although the video diffusion extension in Sec.˜3.2 has introduced spatial-temporal awareness and motion prior from the source video, noticeable illumination and texture flicker persist. To efficiently remove these artifacts, we introduce a two-stage post-optimization framework, as illustrated in parts (b) and (c) of Fig.˜2.
@@ 3.3.1 Stage I: Exposure Alignment
As shown in part (b) of Fig.˜2, the first stage introduces a per-frame appearance embedding $E_{t}$ to compensate for exposure misalignment between adjacent frames. Inspired by kerbl2024hierarchical, we model $E_{t}$ as a $3\times 4$ affine transformation matrix, initialized to the identity and optimized via Adam 2015-kingma. Its supervision combines a photometric term with a flow-warp alignment term using hyperparameter $\lambda_{e}$
@@
Here, $\beta$ is a constant scaling factor, $\xi_{flow}$ and $\xi_{rgb}$ are thresholds set from the statistics of error map $E_{flow}$ and $E_{rgb}$. This soft mask is also applied in the second stage of optimization. As shown in Tab.˜4, soft masking outperforms the hard one in both temporal consistency and prompt alignment.
@@ 3.3.2 Stage II: Optimization over Unique Video Tensor
In the second stage, we refine illumination and texture details. Compared with vanilla video, its canonical representation can incorporate spatial-temporal priors and facilitate consistency ouyang2024codef; shin2024enhancing. But popular NeRF or 3DGS are too complex and costly for learning (cf. Sec.˜2.2). Instead, we compress the video to a one-dimensional RGB vector of shape $(N,3)$, as shown in part (c) of Fig.˜2. Specifically, we define a $d$-dimensional index $\kappa(x,y,t)$ for each pixel based on priors extracted from the source video. An example index could be $[22,127,0,255]$, where the first element is the flow ID (pixels connected by the optical flow predicted by MemFlow share the same flow ID), and the rest are 8-bit quantized RGB values. It is also allowed to extend this 4-element index to more elements with voxel coordinate (from depth projection) or any other cues that indicate spatial–temporal similarity and locality. All pixels with identical $\kappa$ are gathered via averaging to form one element of the one-dimensional vector, where $N$ is the number of unique $\kappa$. Take the source video $\{I_{t}^{in}\}$ as an example, the gathering and scattering operations are formulated as
@@
where $\mathbf{U}$ is referred to as the Unique Video Tensor (UVT). With an appropriate definition of $\kappa$, the scattered $\mathbb{I}_{t}^{in}\left(x,y\right)$ reconstructs the original $I_{t}^{in}\left(x,y\right)$ with minimal information loss, as validated in Tab.˜5. For relighting, the ideal edited video frames must preserve consistent motion and intrinsic image details with the source; thus, they share the same index tensor $\kappa$ for UVT representation. Accordingly, we compress the first-stage output $\tilde{I}_{t}\left(x,y\right)$ into $\mathbf{\tilde{U}}$ via Eq.˜8, which then serves as the primary optimization target. This formulation not only facilitates optimization but also naturally embeds spatial-temporal similarity priors (cf. Sec.˜4.3). With CUDA parallelism, the gathering and scattering process can be performed instantly. The optimization of $\tilde{\mathbf{U}}$ is supervised by
@@
where $\tilde{\mathbb{I}}_{t}\left(x,y\right)=\tilde{\mathbf{U}}\left(\kappa(x,y,t)\right)$, and $\lambda_{tv}$ and $\lambda_{u}\in[0,1]$ balance the loss terms. The total variation loss $\mathcal{L}_{tv}$ suppresses noise. Notably, Eq.˜9 applies SSIM loss instead of photometric loss. This leaves space to fine-grained appearance and illumination adjustment without altering image structure. Finally, the optimized $\widehat{\mathbf{U}}$ is used to reconstruct $\widehat{\mathbb{I}}_{t}(x,y)$ according to Eq.˜8 as the final output.
@@ 4.1 Experiment Setting
Dataset. To comprehensively evaluate the generation capability, we collect video clips with high motion dynamics and broad scene diversity. This subjective evaluation benchmark, as detailed in Tab.˜1, covers scenarios like autonomous driving, robot manipulation, and navigation, as well as drone flight. It includes data from synthetic and realistic environments under various weather conditions. Each clip is a long video with on average 256 frames, making it extremely challenging. To provide a more accurate and robust probe on the performance, we also conduct objective ground-truth-based evaluation on the Virtual KITTI 2 dataset (cabon2020vkitti2). We selected five scenes and relit them to match the illumination of morning, sunset, rain, overcast, and fog settings. Each sequence averages 281 frames at a resolution of 1248×384. To obtain edit prompts, we use some prompts from zhang2025scaling and generate others using COSMOS agarwal2025cosmos.
@@
Metrics. Following prior works qi2023fatezero; li2024vidtome; kara2024rave; wang2025videodirector, we assess the relighting performance along following four dimensions: (i) Temporal consistency is quantified via motion smoothness (Motion-S) huang2024vbench and structural warping error (Warp-SSIM). Motion-S evaluates the continuity and physical plausibility of motion in the edited sequence, whereas Warp-SSIM computes the SSIM between a frame and its warped neighbors using flow from RAFT teed2020raft. (ii) Textual alignment is measured by average CLIP embedding similarity between the text prompt and all edited frames (CLIP-T). (iii) User preference is evaluated by a study on 19 randomly selected videos and 65 valid submissions collected. Participants choose their preferred relighting results among our method and established baselines, from which we derive the Bradley–Terry preference rate (User-PF) bradley1952rank. Additional details are included in the Appendix. (iv) Alignment with Groundtruth is measured using SSIM and LPIPS (kerbl3Dgaussians) with ground truth relighted results. These two metrics replace (ii) and (iii) on Virtual KITTI 2 (cabon2020vkitti2) for a more accurate performance evaluation. (v) Computation efficiency is reported in terms of runtime speed (FPS) and peak GPU memory consumption (VRAM) during editing. All experiments are conducted on a 40GB A100 GPU. Additionally, to appraise the reconstruction quality of UVT, we report average PSNR, SSIM, and LPIPS between the original and reconstructed frames.